OpenStudy (anonymous):

When doing the Thevenin/Norton equivalent circuits, does the resistance somehow get changed into conductance? I'm working on a question and I'm not sure how/why the resistance became conductance.

OpenStudy (kenljw):

The Thevinin and Norton equivatent circuts both have the same internal resistance, usually expressed Rth. Whether some is expressed in conductance, 1/R, is generally irrevalent. When doing Loop/Mesh analysis resistance is expressed in Ohms, in doing Nodal analysis it usually expressed in conductance, 1/Ohms. This allows for straight forward standard Matrixes for solving problems.

OpenStudy (anonymous):

G=1/R = I/V ---** I=GV - G=I/V- V=I/G so if you think, and you wanna introduce the siemens S, I=GV(**).And is also equivalent of (1)S=I/V. if you wanna know something more : http://whatis.techtarget.com/definition/siemens-mho